OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. x + y = 3 2x + 2y = 10

OpenStudy (anonymous):

x+y=3 multiply both sides by 2 so, we have 2x+2y=6 now , it is also given that. 2x+2y=10 therefore, 6=10(false) hence, no solution